# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 8.5 Failed Efforts to Reform the USMA-2006 Mechanism
Despite several attempts with various officials, our efforts to collaborate with the Army on redesigning their branching mechanism were unsuccessful for many years. During this time, my long-time collaborator, Parag Pathak, began incorporating my research on Army branching into his graduate Market Design course at MIT and independently advocated for it to USMA officials.
In 2019, through Pathak, I finally learned why the Army initially chose not to pursue a reform of the USMA-2006 mechanism.
@@
In contrast, the second root cause has more tangible consequences. Some cadets unnecessarily paid the increased price for their assigned branch under the USMA-2006 mechanism, even though they would have received the same branch at its base price without volunteering for BRADSO incentives. This issue was more apparent, as affected cadets could observe that lower OML-ranked cadets received the same branch at the base price. However, this challenge could be managed through a simple manual adjustment by waiving the additional ADSO charge for any affected cadet.
Thus, any anomalous outcomes under the USMA-2006 mechanism were either obscured by its strategy space or could be easily corrected with a simple manual adjustment. Subsequent analyses of cadet preferences found that such failures affected, on average, roughly 2% of cadets under the USMA-2006 mechanism (Greenberg et al., 2024). Much like the benefits of replacing the MCSD mechanism with DA for Turkish college admissions were not deemed substantial enough to justify a costly reform years earlier, the benefits of replacing the USMA-2006 mechanism with the MPCO mechanism were not considered large enough to justify a costly reform of the U.S. Army’s branching system. Regardless of the merits of the alternative, the Army’s mechanism at the time was simply not seen as “bad enough” to necessitate such a reform.
@@ 8.6 ROTC Mechanism
With these specifics in mind, ROTC leadership adopted the ROTC mechanism, which has a strategy space similar to the USMA-2006 mechanism but limits the number of branches cadets can rank. Cadets were allowed to rank up to three branches and indicated which, if any, they were willing to pay an increased price for in exchange for elevated priority in half of the positions. For the allocation process, ROTC, like USMA, used a procedure similar to SSD-OML but applied three criteria for allocation instead of two. Each cadet was evaluated according to their OML standing and placed in the highest-ranked branch for which they qualified, based on the following criteria:
@@
“I just went through the accessions process this year. [$\ldots$] DO NOT GET STUCK IN the 30-40% on the active duty OML. Once in those percentiles it becomes mathematically impossible to get a competitive branch (i.e., infantry and aviation).”
@@
“[$\ldots$] These kids in the DEAD ZONE are 20-year-old cadets being faced with a moral dilemma $\ldots$ do your best and kiss your branch choice goodbye, or screw up and get your branch choice. I think this is a strange choice to put in front of a young cadet.”
To address the failures of the ROTC mechanism, Sönmez (2013) proposed the multi-category cumulative offer (MPCO) mechanism, though in a more flexible manner than Sönmez and Switzer (2013) did for USMA. [^96] Sönmez (2013) refers to the MPCO mechanism as the Cadet Optimal Stable Mechanism (COSM) under Bid for Your Career (BfYC) Priorities.
@@
For USMA, addressing the root causes of the failures was relatively straightforward, as the system did not contain deeply flawed elements tied to its key objectives. This was not the case for ROTC. To remedy certain failures, it was essential to eliminate the dead zones, but this required addressing ROTC’s skill-based distributional requirements through alternative means. I proposed using multiple pricing tiers and potentially increasing the share of flexible-price positions, although other, potentially more effective interventions might also exist. Ultimately, the effectiveness of my proposal remained an empirical question.
@@ 8.7 Talent-Based Branching Program and the USMA-2020 Mechanism
To avoid undermining the TBB program with BRADSO incentives, cadets who volunteered for BRADSO at a branch received higher priority only within their own tier for that branch. Since the USMA-2006 mechanism could not accommodate these changes, a new mechanism was designed and implemented in Fall 2019 for the Class of 2020. We refer to this mechanism, which lasted only one year, as the USMA-2020 mechanism.
Under this new mechanism, the Army retained the same strategy space as in previous years: each cadet submitted a ranking of branches and identified any branches for which they were willing to pay an additional price to receive elevated priority in a subset of positions. Given a profile of cadet strategies, the Army constructed an adjusted priority order for each branch, incorporating both TBB ratings and willingness to pay the increased price. The USMA-2020 mechanism then determined branch assignments using Gale and Shapley’s individual-proposing Deferred Acceptance algorithm, based on the submitted preferences and adjusted priorities.
@@
This was a somewhat rushed design. In addition to inheriting the two root causes underlying the failures of the USMA-2006 mechanism, the USMA-2020 mechanism faced a third vulnerability: Even though the number of positions allocated via the BRADSO program was kept at 25% of the total capacity at each branch, cadets received elevated priority for “all” positions in a branch upon volunteering for BRADSO incentives. While this design choice allowed the heterogeneity in branch priorities to be accommodated using the vanilla version of the Deferred Acceptance algorithm, [^97] Indeed, the following quote from the U.S. Army News story O’Connor (2019) suggests that the design of the USMA-2020 mechanism was inspired by the NRMP’s medical match: “The cadets’ branch rankings and the branches’ cadet preferences will then determine a cadet’s branch using a modified version of the National Resident Matching Program’s algorithm, which won a Nobel Prize for Economics in 2012 and pairs medical school graduates with residency programs.” it also resulted in a more problematic version of the no priority reversal failures that could not be manually corrected.
@@ 8.8 Multi-Price Cumulative Offer Mechanism as a Minimalist Reform
Over the years, the mission of the Army’s branching system has evolved. Initially, reflecting the Army’s emphasis on hierarchy, the system’s primary role was to allocate positions based solely on merit. With the introduction of the BRADSO incentives program in the mid-2000s, a second objective emerged: increasing officer retention. Later, the TBB program introduced a third goal—improving the match between talent and specialties.
Maintaining trust in the system was another key objective, a task that was relatively straightforward when merit-based allocation was the sole focus but became increasingly challenging as additional goals were introduced.
In light of these evolving objectives, this section will explore how axiomatic methodology is particularly well suited to addressing these challenges and why the MPCO mechanism is uniquely compelling for the Army’s mission.
@@
Over the years, this axiom—first formalized in Balinski and Sönmez (1999)—and its generalizations have been central to a range of real-world reforms, contributing significantly to several “discovery-invention cycles.” The Army’s branching mechanism was the second of these applications, following school choice.
The essence of the no justified envy axiom is to ensure that position assignments respect the “property rights” of individuals as defined by institutions. The original version of this axiom, presented in Definition 5, applies to settings where all positions within an institution are allocated based on a single priority ranking. While the no priority reversal axiom, outlined in Definition 10, is a direct extension of the no justified envy principle and central to the Army’s mission, it addresses cadet objections to branch assignments of others only if they are willing to retain their prices. However, this alone does not fully capture the full scope of the no justified envy axiom in the current setting.
@@ 8.8.1 Price-responsiveness Schemes and Respecting Them
What, then, is missing from the no priority reversal axiom in terms of fully respecting property rights established through two criteria? To explore this, let’s consider an example based on the scoring-based price-responsiveness scheme, inspired by the scenario in Section 4.1.989898Interestingly, according to Zhou and Wang (2024), scoring-based price responsiveness was used in several major Chinese cities, including Beijing, Shanghai, and Tianjin, from the 1990s until mid-2010s for high school admissions. Under this policy, students could receive fixed score boosts for a fraction of the seats—often across multiple tiers in some cities—in exchange for paying higher tuition levels.
@@
Consider a school with 50 seats and a base tuition of \$10,000. For 40 of the seats, applicant priority is based on a standardized test score out of 100. For the remaining 10 seats, priority is determined by an adjusted score, where applicants who pay an increased tuition of \$25,000 gain an extra 20 points added to their test score.
@@
For Brian to have a valid claim to Anna’s seat by offering the increased tuition, fewer than 10 seats must have been awarded at the higher price. Let’s assume this is the case. Additionally, Brian must have a score higher than 45, since paying the extra tuition only gives him 20 additional points. Therefore, if his score is 50, he has a valid claim to Anna’s seat at the increased tuition. However, if his score is 40, he does not, even if fewer than 10 flexible-price seats were awarded at the increased price.
@@
Definition 16 includes an additional condition compared to 15 because, while awarding a position by reducing its price is always feasible, awarding a position by increasing its price may be infeasible due to the cap on the number of flexible-price positions.
@@ 8.8.2 A Characterization of the MPCO Mechanism
Starting with school choice reforms in Boston, New York City, Chicago, and England in the mid-2000s and early 2010s, DA emerged as one of the leading student assignment mechanisms worldwide. Earlier, in Section 6.4, Theorem 8 established that DA is the unique mechanism that satisfies individual rationality, non-wastefulness, no justified envy, and strategy-proofness. Given the prominent role played by the no justified envy and strategy-proofness axioms in these reforms, this result naturally contributed to the widespread acceptance of DA.
@@
The multi-price cumulative offer mechanism is the only mechanism that satisfies individual rationality, non-wastefulness, no justified envy, and strategy-proofness.
@@
“Tuition for public high schools is based on the student’s exam score. Since scores are the only admission criteria, schools set a cutoff for normally admitted students. Normal students pay an annual tuition of 1600 yuan (roughly \$260 in 2013), ensuring affordability. ZX students’ tuition depends on their score: scores within 10 points of the cutoff result in a $3$.3 yuan annual fee; scores 11–20 points above the cutoff pay 5000 yuan; and scores 21–30 points above the cutoff pay 6000 yuan. No school can admit a ZX student with a score more than 30 points below its cutoff.” (p. 1157)
@@
Beginning with the branching of the Class of 2021 cadets to Army branches in Fall 2020, our collaboration with the Army led to the adoption of the MPCO mechanism at USMA, incorporating the Army’s dual-criteria priority structure (Greenberg et al., 2024). Impressed by the success of the new system, the Army decided to extend it to ROTC that same year.
@@
9 Design of Property Rights
In Sections 6 and 8, I argue that the failure to satisfy the axioms of no justified envy and strategy-proofness has been a major catalyst for reforms in both school choice and the Army’s branching system over the past two decades. Consequently, many school choice and college admissions systems, as well as the U.S. Army, have moved away from mechanisms that violate these principles. Instead, they have adopted the DA mechanism or its generalized version, the cumulative offer mechanism, primarily because these mechanisms satisfy the desired criteria. Indeed, as established in Theorems 8 and 10, in many settings, these are the only mechanisms that fulfill these principles along with the basic requirements of individual rationality and non-wastefulness.
@@ 9 Design of Property Rights
Provided that each institution’s choice rule satisfies certain technical conditions, mechanisms like DA or the cumulative offer mechanism can be understood as instruments that extend the implementation of “property rights” from single-institution settings to multi-institution settings. Consequently, exploring the technical conditions that ensure this transition has been an active area of matching theory since the seminal work of Hatfield and Milgrom (2005). [^99] Technical conditions such as variants of the substitutability condition and the IIC condition ensure that the algorithms of these mechanisms not only terminate but also generate outcomes that satisfy the no justified envy axiom, thereby enforcing the underlying property rights in multi-institution settings. Additionally, LAC condition ensures that these mechanisms satisfy the strategy-proofness axiom, thus enforcing these property rights in an incentive-compatible manner. See Hatfield et al. (2021) for the mildest known conditions that serve these objectives.
Despite the theoretical interest in choice rules, our use-inspired research had never interfered with institutional choice rules—unless they were among the root causes of failures in the underlying mechanisms—up until the early 2010s. The minimalist mindset that guided our holistic approach to both research and policy was a key reason for this stance. This remained true until Boston Mayor Thomas Menino’s pivotal 2012 State of the City Address, where he pledged to increase the fraction of students assigned to schools in their neighborhoods.100100100One instance where a choice rule directly contributed to failures was the ROTC branching mechanism, leading to a proposed mechanism with an alternative choice rule, as discussed in Section 8.6. From a pragmatic standpoint, interfering with choice rules would not only deviate from minimalist market design but also risk alienating interest groups negatively impacted by our proposals.
This stance was also influenced by the implicit assumption we had made up to that point: that the selection of choice rules by institutions reflected their intended policies. Thus, unless it was a technical necessity, we saw no reason to propose changes that might interfere with these policies. However, following Menino’s State of the City Address, it became clear that this may not always be the case, and that designing choice rules that accurately reflect an institution’s intended policies may itself be an essential exercise in market design (Dur et al., 2018b).

### v4

@@ 8.5 Failed Efforts to Reform the USMA-2006 Mechanism
Despite several attempts with various officials, our efforts to collaborate with the Army on redesigning their branching mechanism were unsuccessful for many years. During this time, my long-time collaborator Parag Pathak began incorporating my research on Army branching into his graduate Market Design course at MIT and independently advocated for it to USMA officials.
In 2019, through Pathak, I finally learned why the Army initially chose not to pursue a reform of the USMA-2006 mechanism. I later came to interpret this explanation as the Army’s official justification for its decision. In retrospect, this is unsurprising: the Army was unlikely to adopt unsolicited reform proposals from outside researchers without a strong internal rationale and its own process of institutional vetting.
@@
By contrast, the second root cause had more tangible consequences. Some cadets unnecessarily paid the increased price for their assigned branch under the USMA-2006 mechanism, even though they would have received the same branch at its base price without volunteering for BRADSO incentives. This issue was more apparent, as affected cadets could observe that lower OML-ranked cadets received the same branch at the base price. However, this challenge could be managed through a simple manual adjustment by waiving the additional ADSO charge for any affected cadet.
Thus, anomalous outcomes under the USMA-2006 mechanism were either obscured by its strategy space or could be corrected manually, and they affected only about 2% of cadets on average (Greenberg et al., 2024). Much like the case of Turkey, where the benefits of replacing MCSD with DA were not deemed large enough to justify reform, the benefits of replacing the USMA-2006 mechanism with the MPCO mechanism were likewise judged insufficient. Regardless of the merits of the alternative, the Army’s mechanism at the time was simply not seen as “bad enough” to prompt change.
@@ 8.6 ROTC Mechanism
With these specifics in mind, ROTC leadership adopted a mechanism that shares the strategy space of the USMA-2006 mechanism but limits the number of branches cadets may rank.
ROTC Mechanism.
Cadets could rank up to three branches and indicated which, if any, they were willing to pay an increased price for in exchange for elevated priority in half of the positions. For the allocation process, ROTC, like USMA, used a procedure similar to SSD-OML but applied three criteria for allocation instead of two. Each cadet was evaluated according to their OML standing and placed in the highest-ranked branch for which they qualified, based on the following criteria:
@@
Figure 23: ROTC branch assignment results for the year 2011. For each branch, the blue region represents the portion of the OML where cadets secured a position within the top 50 percent of positions at the base price. The purple region corresponds to cadets who obtained a position between the 50th and 65th percentiles at the increased price, while the brown region indicates cadets who secured a slot in the last 35 percent, also at the increased price. The numbers in these regions represent the number of cadets in each bracket. The blank space between the purple and light brown regions marks the “dead zone” for the eight most popular branches: Aviation (AV), Infantry (IN), Medical Service (MS), Armor (AR), Engineering (EN), Military Intelligence (MI), Military Police (MP), and Financial Management (FI). The less sought-after branches, which do not have a dead zone, include Signal Corps (SC), Engineering reserves for those with engineering degrees (EN(D)), Adjutant General’s Corps (AG), Field Artillery (FA), Air Defense Artillery (AD), Transportation Corps (TC), Quartermaster Corps (QM), Ordnance Corps (OD), and Chemical Corps (CM). Figure reproduced from Figure 1 in Sönmez (2013), Journal of Political Economy, https://doi.org/10.1086/669915, originally published in grayscale. Reprinted with permission. © 2013 by The University of Chicago.
@@
“I just went through the accessions process this year. $\ldots$] DO NOT GET STUCK IN the 30–40% on the active duty OML. Once in those percentiles it becomes mathematically impossible to get a competitive branch (i.e., infantry and aviation).”
@@
$\ldots$] These kids in the DEAD ZONE are 20-year-old cadets being faced with a moral dilemma $\ldots$ do your best and kiss your branch choice goodbye, or screw up and get your branch choice. I think this is a strange choice to put in front of a young cadet.”
To address the failures of the ROTC mechanism, Sönmez (2013) proposed the multi-category cumulative offer (MPCO) mechanism, though in a more flexible manner than Sönmez and Switzer (2013) did for USMA.109109109Sönmez (2013) refers to the MPCO mechanism as the Cadet Optimal Stable Mechanism (COSM) under Bid for Your Career (BfYC) Priorities.
@@
For USMA, addressing the root causes of the failures was relatively straightforward, as the system did not contain deeply flawed elements tied to its key objectives. This was not the case for ROTC. To remedy certain failures, it was essential to eliminate the dead zones, but this required addressing ROTC’s skill-based distributional requirements through alternative means. I proposed using multiple pricing tiers and potentially increasing the share of flexible-price positions, although other, potentially more effective interventions might also exist. Ultimately, the effectiveness of my proposal remains an empirical question.
@@ 8.7 Talent-Based Branching Program and the USMA-2020 Mechanism
To avoid undermining the TBB program with BRADSO incentives, cadets who volunteered for BRADSO at a branch received higher priority only within their own tier for that branch. Since the USMA-2006 mechanism could not accommodate these changes, a new mechanism was designed and implemented in Fall 2019 for the Class of 2020.
USMA-2020 Mechanism.
Under this new mechanism, which lasted only one year, the Army retained the same strategy space as in previous years: each cadet submitted a ranking of branches and identified any branches for which they were willing to pay an additional price in exchange for elevated priority in a subset of positions. Given a profile of cadet strategies, the Army constructed an adjusted priority order for each branch, incorporating both TBB ratings and willingness to pay the increased price. The USMA-2020 mechanism then determined branch assignments using Gale and Shapley’s individual-proposing deferred acceptance algorithm, based on the submitted preferences and adjusted priorities.
@@
This was a somewhat rushed design. In addition to inheriting the two root causes underlying the failures of the USMA-2006 mechanism, the USMA-2020 mechanism faced a third vulnerability: Even though the number of positions allocated via the BRADSO program was kept at 25% of the total capacity at each branch, cadets received elevated priority for “all” positions in a branch upon volunteering for BRADSO incentives. While this design choice allowed the heterogeneity in branch priorities to be accommodated using the vanilla version of the deferred acceptance algorithm,110110110Indeed, the following quote from the U.S. Army News story O’Connor (2019) suggests that the design of the USMA-2020 mechanism was inspired by the NRMP’s medical match: “The cadets’ branch rankings and the branches’ cadet preferences will then determine a cadet’s branch using a modified version of the National Resident Matching Program’s algorithm, which won a Nobel Prize for Economics in 2012 and pairs medical school graduates with residency programs.” it also resulted in a more problematic version of the no priority reversal failures that could not be manually corrected.
@@ 8.8 Multi-Price Cumulative Offer Mechanism as a Minimalist Reform
Over the years, the mission of the Army’s branching system has evolved. Initially, reflecting the Army’s emphasis on hierarchy, the system’s primary role was to allocate positions based solely on merit. With the introduction of the BRADSO incentives program in the mid-2000s, a second objective emerged: improving officer retention. Later, the TBB program added a third goal—enhancing the match between talent and specialties.
Maintaining trust in the system became another key objective, a task that was relatively straightforward when merit-based allocation was the sole focus but increasingly challenging as additional goals were introduced.
In light of these evolving objectives, this section examines why axiomatic methodology is particularly well suited to addressing these challenges and why the MPCO mechanism is uniquely compelling for the Army’s mission.
@@
Over the years, this axiom—first formalized in Balinski and Sönmez (1999)—and its generalizations have been central to a range of real-world reforms, contributing significantly to several “discovery–invention cycles” illustrated in Figure 36. The Army’s branching mechanism was the second of these applications, following school choice.
The essence of the no justified envy axiom is to ensure that position assignments respect the “property rights” of individuals as defined by institutions. The original version of this axiom, presented in Definition 5, applies to settings where all positions within an institution are allocated based on a single priority ranking. While the no priority reversal axiom, given in Definition 10, is a direct extension of the no justified envy principle and central to the Army’s mission, it addresses cadet objections to branch assignments of others only if they are willing to retain their prices. However, this alone does not fully capture the full scope of the no justified envy axiom in the current setting.
@@ 8.8.1 Price-responsiveness Schemes and Respecting Them
What, then, is missing from the no priority reversal axiom in terms of fully respecting property rights established through two criteria? To explore this, let’s consider an example based on the scoring-based price-responsiveness scheme, inspired by the scenario in Section 4.1.
@@
Consider a school with 50 seats and a base tuition of \$10,000. For 45 of the seats, applicant priority is based on a standardized test score out of 100. For the remaining 5 seats, priority is determined by an adjusted score, where applicants who pay an increased tuition of \$25,000 gain an extra 20 points added to their test score.
@@
For Brian to have a valid claim to Anna’s seat by offering the increased tuition, fewer than 5 seats must have been awarded at the higher price. Let’s assume this is the case. Additionally, Brian must have a score higher than 45, since paying the extra tuition only gives him 20 additional points. Therefore, if his score is 50, he has a valid claim to Anna’s seat at the increased tuition. However, if his score is 40, he does not, even if fewer than 5 flexible-price seats were awarded at the increased price. (See Figure 24 for an illustration.)
Figure 24: Legitimate claim for a price-reduced version of another individual’s assignment. The school’s base tuition is \$10K, and seats are allocated based on applicants’ scores on an entry test. For five seats, applicants may obtain a 20-point score boost by paying an additional \$15K, for a total tuition of \$25K. Suppose that, under the current allocation, fewer than five students pay the higher tuition of \$25K. Anna secures a seat at the base tuition of \$10K with a score of 65, while Brian, with a score of 50, remains unassigned. Because Brian’s score is lower than Anna’s, he cannot contest her seat at the base tuition (indicated by a crossed-out dotted arrow). However, since not all five flexible-price seats are filled at the higher tuition of \$25K, he may contest Anna’s assignment (indicated with a bold arrow) if he is willing to pay the increased tuition.
@@
Definition 16.
  1. 1.
  2. 2.
Definition 16 omits the last condition in Definition 15 because, whereas awarding a position at a higher price may be infeasible due to the cap on flexible-price positions, awarding a position at a lower price is always feasible.
@@ 8.8.2 A Characterization of the MPCO Mechanism
Starting with school choice reforms in Boston, New York City, Chicago, and England in the mid-2000s and early 2010s, the Deferred Acceptance (DA) mechanism emerged as one of the leading student assignment mechanisms worldwide. Earlier, in Section 6.4, Theorem 8 established that DA is the unique mechanism that satisfies individual rationality, non-wastefulness, no justified envy, and strategy-proofness. Given the prominent role played by the no justified envy and strategy-proofness axioms in these reforms, this result naturally contributed to the widespread acceptance of DA.
@@
With this adjustment in place, Theorem 8 extends directly:
@@
The multi-price cumulative offer mechanism is the unique mechanism that satisfies individual rationality, non-wastefulness, no justified envy, and strategy-proofness.
@@
“Tuition for public high schools is based on the student’s exam score. Since scores are the only admission criteria, schools set a cutoff for normally admitted students. Normal students pay an annual tuition of 1600 yuan (roughly \$260 in 2013), ensuring affordability. ZX students’ tuition depends on their score: scores within 10 points of the cutoff result in a 3333.3 yuan annual fee; scores 11–20 points above the cutoff pay 5000 yuan; and scores 21–30 points above the cutoff pay 6000 yuan. No school can admit a ZX student with a score more than 30 points below its cutoff.” (p. 1157)
@@
Beginning with the branching of the Class of 2021 cadets to Army branches in Fall 2020, our collaboration with the Army led to the adoption of the MPCO mechanism at USMA, incorporating the Army’s dual-criteria priority structure (Greenberg et al., 2024). Encouraged by the success of the new system, the Army decided to extend it to ROTC that same year.
@@
9 Reserve Systems
In Sections 6 and 8, I argue that the failure to satisfy the axioms of no justified envy and strategy-proofness has been a major catalyst for reforms in both school choice and the Army’s branching system over the past two decades. As a result, many school choice and college admissions systems, as well as the U.S. Army, have abandoned mechanisms that violate these desiderata. They have instead adopted the DA mechanism or its generalized version, the cumulative offer mechanism, primarily because these mechanisms satisfy them. As established in Theorems 8 and 10, in many settings these are, in fact, the only mechanisms that meet these desiderata together with the basic axioms of individual rationality and non-wastefulness.
@@ 9 Reserve Systems
Provided that each institution’s choice rule satisfies certain technical conditions, mechanisms like DA or the cumulative offer mechanism can be understood as instruments that extend the implementation of “property rights” from single-institution settings to multi-institution settings. Consequently, exploring the technical conditions that ensure this transition has been an active area of matching theory since the seminal work of Hatfield and Milgrom (2005).111111111Technical conditions such as variants of the substitutability condition and the IIC condition ensure that the algorithms of these mechanisms not only terminate but also generate outcomes that satisfy the no justified envy axiom, thereby enforcing the underlying property rights in multi-institution settings. Additionally, LAC condition ensures that these mechanisms satisfy the strategy-proofness axiom, thus enforcing these property rights in an incentive-compatible manner. See Hatfield et al. (2021) for the mildest known conditions that serve these objectives.
Up until the early 2010s, despite the theoretical interest in choice rules, our practical design efforts had never interfered with institutional choice rules, unless they were among the root causes of failures in the underlying mechanisms. The minimalist mindset guiding our holistic approach to both research and policy was a key reason for this stance. This remained true until Boston Mayor Thomas Menino’s pivotal 2012 State of the City Address, in which he pledged to increase the fraction of students assigned to schools in their neighborhoods.112112112One instance where a choice rule directly contributed to failures was the ROTC branching mechanism, leading to a proposed mechanism with an alternative choice rule, as discussed in Section 8.6. From a pragmatic standpoint, interfering with choice rules would not only deviate from minimalist market design but also risk alienating interest groups adversely affected by our proposals.
This stance was also shaped by an implicit assumption we had held up to that point: that institutions’ choice rules faithfully reflected their intended policies. Accordingly, unless technical necessity dictated otherwise, we saw no reason to propose changes that might interfere with those policies. Menino’s State of the City Address, however, made clear that this need not always be the case, and that designing choice rules to align with an institution’s actual policy objectives may itself constitute an essential exercise in market design (Dur et al., 2018b).